# Evaluating Low-Light Image Enhancement Across Multiple Intensity Levels

### v1

@@ Abstract
Figure 1: Illustration of our capture setup. A set of controllable lights illuminates the scene. For each scene, we capture 11 images by varying the light intensity from minimum to maximum at 10% intervals (levels). Camera parameters (aperture, exposure time, ISO) remain fixed, and images are captured in unprocessed RAW format. The image captured at maximum intensity serves as the ground truth, while all other images serve as low-light inputs for training, validation, and testing. Scenes are captured with both a DSLR and a smartphone, and scene illuminance is measured with a photometer.
@@ 1 Introduction
Images taken in low-light environments are corrupted by sensor noise and diminished color saturation. Simple digital exposure adjustments, such as scaling the image’s digital values, result in poor image quality due to high levels of sensor noise. Consequently, deep learning techniques have been developed to directly enhance low-light images, efficiently reducing noise and improving color and texture (e.g., [RetinexNet, KinD, Ruas, Retinexformer, LLFormer]). The success of these approaches is heavily dependent on how the training data is collected.
@@
Using the MILL dataset, we analyze how current state-of-the-art methods perform under varying lighting conditions and find that model performance varies significantly across different intensity ranges. Based on our findings, we propose an improvement over the best-performing method, Retinexformer [Retinexformer]. We propose to disentangle scene and illumination information in the network’s latent features by leveraging the multi-level nature of our dataset. We demonstrate that our simple modification improves PSNR by $10$ dB for the DSLR and $2$ dB for the smartphone camera on Full HD images.
@@ 2.1 Low-Light Datasets
Early LLIE datasets, such as VV [VV] and LIME [LIME], contained only unpaired low-light images (15 and 10 samples, respectively) without corresponding well-exposed references. For this reason, Wei et al. [Chen2018Retinex] introduced the LOw Light paired dataset (LOL) to allow end-to-end training of LLIE models. The LOL dataset proved valuable to the research community, enabling end-to-end training of methods. LoLv1 contains images captured under different camera settings to capture the same scene under low-light and well-lit conditions. The LOLv1 dataset consists of $500$ images, of which $485$ are for training and $15$ are for testing. An extension of the LoLv1 dataset, LoLv2, was later introduced [lolv2]. In LoLv2, the authors introduced two variants: one following the LoLv1 methodology and the other generating the low-light image synthetically from the well-lit counterpart. LoLv2 contains $689$ training scenes and $100$ test scenes. A major issue with these datasets is that they contain images from the same scene in both the training and test sets, potentially affecting generalization.
Several datasets were subsequently introduced to address the limitations of early LLIE benchmarks, primarily in terms of scale and diversity. The DPED [DPED] dataset provided images captured across multiple smartphone cameras, enabling cross-device evaluation. Deep-UPE [DEEPUPE] emphasized extremely low-light scenarios with more challenging exposure conditions. The LSRW [LRSW] dataset expanded camera diversity by including both DSLR and smartphone captures, recognizing the distinct image formation characteristics of different sensor types. More recently, LLIV-Phone [LLIV-Phone] introduced temporal information by capturing video sequences under low-light conditions, allowing methods to exploit inter-frame correlations. This video-based approach was further explored in the DID [DID] and SDSD [SDSD] datasets, which provided paired low-light and normal-light video sequences for dynamic scene enhancement.
A significant methodological shift came with the SID [SID] dataset, which captured image pairs in RAW format rather than processed RGB, enabling methods to leverage the complete sensor information before in-camera processing. To scale dataset creation, VE-LOL [VE-LOL] adopted a synthetic approach, darkening well-exposed images and adding synthetic noise patterns to simulate sensor characteristics at high ISO settings. Recently, the BVI-LowLight dataset [BVI-lowlight] was introduced, containing 40,000 images of objects captured at different ISO settings. Additionally, to obtain more training data, some LLIE methods use exposure-correction datasets such as PHOS [Phos] and SICE [SICE]. Despite these advances in scale, sensor diversity, temporal modeling, and data modality, existing datasets either capture each scene under a single low-light condition or rely on modifying the camera parameters.
@@ 2.2 Low-Light Image Enhancement Methods
Early LLIE methods built upon the seminal Retinex algorithm [land1964retinex], which decomposed images into reflectance and illumination components, inspiring variants including Multi-Scale Retinex [rahman1996multi], SRIE [fu2016weighted], and Milano-Retinex [milano1]. Methods such as LIME [LIME] and NPE [wang2013naturalness] demonstrated strong performance by leveraging natural image statistics without training data. However, these traditional approaches have been largely superseded by deep learning methods.
Early end-to-end deep learning methods include SID [SID] for RAW images and RetinexNet [Chen2018Retinex] for RGB inputs. Subsequent approaches introduced various architectural innovations: GLADNet [wang2018gladnet] combined global illumination and local detail modules; KinD [KinD] and KinD++ [zhang2021beyond] adopted Retinex-inspired decomposition strategies; and Yang et al. [yang2020fidelity] incorporated adversarial learning. Recent methods leverage transformer architectures (LLFormer [LLFormer, LLFormer2], Retinexformer [Retinexformer]) and diffusion models (Diff-Retinex [yi2023diff], PyDiff [pydiff]). Alternative formulations include specialized color spaces [hvi] and pixel-wise mean estimation losses [liao2025gt]. Some methods jointly address enhancement and degradation removal, such as DarkIR [darkir] and LEDNet [zhou2022lednet]. To avoid paired training data requirements, unsupervised approaches have been proposed, including EnlightenGAN [jiang2021enlightengan], Zero-DCE [guo2020zero], SCI [SCI], and lightweight RUAS [Ruas].
However, all existing methods have been evaluated exclusively on fixed single low-light inputs without considering behavior across varying illumination levels. We benchmark state-of-the-art LLIE methods across different brightness conditions and propose two simple modifications to Retinexformer [Retinexformer] that leverage our multi-illumination dataset to improve robustness across intensity levels.
@@ 3 Multi-Illumination Low-Light (MILL) Dataset
Figure 3: Example scenes from our dataset at different levels for both the DSLR camera (first and second rows) and the smartphone camera (third row). We show the three lowest values to illustrate how intensity varies across successive levels. We also display higher levels to show they remain noticeably underexposed compared to the ground truth.
Existing LLIE datasets present two critical limitations: (1) they either contain a single severely underexposed image per scene (e.g. [lolv2, Chen2018Retinex]) or (2) simulate brightness variations via camera parameter adjustments or post-processing (e.g. [BVI-lowlight, SICE, Phos]). This constraint limits real-world applicability, where low-light conditions span a continuous range of intensities.
To quantify this limitation, we simulated varying brightness levels on the LoLv1 dataset [RetinexNet] by blending input images with their ground truth counterparts at different ratios. This blending reduces degradation severity by simulating intermediate brightness levels, theoretically making enhancement easier. As shown in Table 1, both Retinexformer [Retinexformer] and CIDNet [hvi] performed worse on the blended versions (with 0.2 and 0.5 ground truth mixing ratios) than on the original dataset, as measured by $\Delta E_{76}$ and PSNRL. This result occurs because models trained on fixed brightness levels fail to generalize across different intensities. Figure 2 illustrates this problem. Oversaturation in the output increases proportionally with input brightness, a clear evidence that intermediate brightness levels are absent from training data. This lack of intensity diversity severely limits the practical applicability of LLIE methods, as real-world deployment requires robustness across varying brightness conditions. To study this problem and address this limitation, we introduce a novel LLIE dataset featuring multiple brightness levels per scene with fixed camera parameters, enabling more robust training, benchmarking, and evaluation of LLIE methods across multiple intensity levels.
Our primary objective is to capture a high-quality, well-calibrated dataset for evaluating and training LLIE methods across different intensity levels. To ensure consistency and eliminate uncontrollable variables, we captured all images in a controlled indoor environment without windows or external light sources. We used a dedicated room with a platform for placing different floor backgrounds and objects, equipped with programmable lighting to precisely control brightness levels. Images were captured using two devices: a Nikon D5200 DSLR camera and a Samsung Galaxy S7 smartphone. The smartphone provides a contrasting capture profile compared to the DSLR, enabling evaluation across different sensor characteristics. Figure 1 provides a schematic overview.
@@
RUAS [Ruas] 0.003 CVPR’21 25.46 16.63 45.33 9.86 57.47 6.93 62.99 5.86 67.14 5.18
LLFormer [LLFormer2] 24.52 AAAI’23 16.37 20.88 13.73 21.79 13.34 22.06 13.17 22.25 12.90 22.55
KinD [KinD] 1.20 ACMM’19 23.88 16.71 17.62 21.79 14.87 21.34 14.49 21.63 15.00 21.36
FourLLIE [wang2023fourllie] 0.12 ACMM’23 24.51 17.29 22.79 17.66 26.64 14.97 28.79 14.07 30.79 13.35
SCI [SCI] 0.0003 CVPR’22 24.05 16.02 17.99 21.42 25.89 15.69 31.66 13.24 38.38 11.23
@@ 4 Method using New Loss Terms
Most current LLIE architectures follow a UNet-like structure, comprising an encoder and a decoder. We aim to disentangle the latent features extracted by the architecture’s bottleneck. We adopt Retinexformer [Retinexformer] as our baseline architecture due to its strong performance in our benchmark evaluation (see Section 5.1).
@@ 4.1 Intensity Prediction Loss
KinD [KinD] 20.409 18.157 0.779 0.219 16.782 1.871 4.383 15.948
FourLLIE [wang2023fourllie] 15.486 13.534 0.687 0.241 26.759 3.514 5.398 19.201
SCI [SCI] 15.329 13.006 0.609 0.270 28.669 3.407 5.791 26.788
@@
Smartphone Unprocessed 19,155 17,573 0,511 0,215 17,832 2.718 5,088 20,895
RUAS [Ruas] 6,920 5,082 0,267 0,722 62,707 6.246 10,556 71,760
@@
KinD [KinD] 19,462 17,452 0,536 0,259 17,252 1.980 3,546 22,950
FourLLIE [wang2023fourllie] 21,377 18,510 0,539 0,236 16,976 3.279 4,831 22,159
SCI [SCI] 15,960 13,208 0,437 0,330 29,477 3.314 5,475 24,372
MirNet [mirnet] 20,760 18,100 0,614 0,219 15,231 1.720 3,537 23,269
@@ 5.1 Benchmark on MILL
We benchmark mainstream LLIE methods by retraining them on our MILL-s dataset using their officially released code. Our evaluation includes unsupervised methods (RUAS [Ruas], SCI [SCI]), a Retinex-based approach (KinD [KinD]), transformer-based methods built on Restormer [restormer] (LLFormer [LLFormer], Retinexformer [Retinexformer], PromptNorm [promptnorm]), a frequency-domain method (FourLLIE [wang2023fourllie]), image restoration approaches (MIRNet [mirnet], DarkIR [darkir]), and specialized LLIE methods (HVI-CIDNet [hvi], GT-Mean [liao2025gt]). We also evaluate our proposed modifications to Retinexformer.
Input SCI [SCI] GT-Mean [liao2025gt] PromptNorm [promptnorm] Retinexformer [Retinexformer] Ours Ground Truth
Figure 4: Visual comparison on MILL-s. From left to right: input, SCI [SCI], GT-Mean [liao2025gt], PromptNorm [promptnorm], Retinexformer [Retinexformer], Ours, and ground truth. We show three examples with zoomed-in regions below each. First two images: DSLR; last image: smartphone.
@@
Figure 6: Outdoor examples from the DICM [DICM] (first row) and SICE [SICE] (second row) of Retinexformer trained on LoLv1, our baseline with the two proposed additional loss terms independently, and our final approach.
@@
Table 3 reports performance averaged across all intensity levels for both the DSLR and smartphone splits of MILL-s, using three full-reference metrics: PSNRL, PSNRC (on RGB), and SSIM; one perceptual full-reference metric: LPIPS [lpips]; two color similarity metrics: $\Delta E_{76}$ and MS-SWD [ms-swd]; and two non-reference metrics: NIQE [niqe], and Brisque [brisque]. Surprisingly, several methods fail to improve upon the input images on average. As demonstrated in Table 2, these methods enhance extremely low-light images but degrade moderately underexposed images, resulting in net negative impact. This observation highlights the difficulty of achieving robust LLIE across variable intensity levels. Retinexformer achieves the best performance, followed closely by PromptNorm and LLFormer. This indicates that Restormer-based architectures perform well for this task. We therefore select Retinexformer as our baseline due to its superior performance and parameter efficiency. Our proposed method further improves upon Retinexformer by 4.6dB in PSNRL, 3.8dB in PSNRC, and 1.2 in $\Delta E_{76}$ on the DSLR split and 0.5dB in PSNRL, PSNRC, and 0.5 $\Delta E_{76}$ units on the smartphone split. The smartphone split proves more challenging across all methods due to inferior sensor quality. However, our proposed auxiliary loss terms achieve superior performance in most metrics.
@@ 5.3 FullHD Experiments and Ablation
Our proposed modifications outperform the baseline model across all metrics. On the DSLR split, we observe improvements of approximately 10 dB in PSNRL and PSNRCC, 0.03 in SSIM, and 5 in $\Delta E{76}$. The smartphone split exhibits smaller but consistent gains due to sensor limitations; nevertheless, our method maintains a clear performance advantage over the baseline. Notably, the intensity prediction loss yields larger improvements than the scene content loss when applied independently. However, combining both losses delivers the strongest performance, as effective feature disentanglement requires their joint optimization.

### v2

@@ 1 Introduction
Figure 1: Illustration of our capture setup. A set of controllable lights illuminates the scene. For each scene, we capture $1$ images by varying the light intensity from minimum to maximum at 10% intervals (levels). Camera parameters (aperture, exposure time, ISO) remain fixed, and images are captured in unprocessed RAW format. The image captured at maximum intensity serves as the ground truth, while all other images serve as low-light inputs for training, validation, and testing. Scenes are captured with both a DSLR and a smartphone, and scene illuminance is measured with a photometer.
Images taken in low-light environments are corrupted by sensor noise and diminished color saturation. Simple digital exposure adjustments, such as scaling the image’s digital values, result in poor image quality due to high levels of sensor noise. Consequently, deep learning techniques have been developed to directly enhance low-light images, efficiently reducing noise and improving color and texture (e.g., [33, 47, 23, 2, 43]). The success of these approaches is heavily dependent on how the training data is collected.
@@
Using the MILL dataset, we analyze how current state-of-the-art methods perform under varying lighting conditions and find that model performance varies significantly across different intensity ranges. Based on our findings, we propose an improvement over the best-performing method, Retinexformer [2]. We propose to disentangle scene and illumination information in the network’s latent features by leveraging the multi-level nature of our dataset. We demonstrate that our simple modification improves PSNR by $10$ dB for the DSLR and $2$ dB for the smartphone camera on Full HD images.
@@ 2.1 Low-Light Datasets
Early LLIE datasets, such as VV [27] and LIME [8], contained only unpaired low-light images (15 and 10 samples, respectively) without corresponding well-exposed references. For this reason, Wei et al. [35] introduced the LOw Light paired dataset (LOL) to allow end-to-end training of LLIE models. The LOL dataset proved valuable to the research community, enabling end-to-end training of methods. LoLv1 contains images captured under different camera settings to capture the same scene under low-light and well-lit conditions. The LOLv1 dataset consists of $500$ images, of which $485$ are for training and $15$ are for testing. An extension of the LoLv1 dataset, LoLv2, was later introduced [39]. In LoLv2, the authors introduced two variants: one following the LoLv1 methodology and the other generating the low-light image synthetically from the well-lit counterpart. LoLv2 contains $689$ training scenes and $100$ test scenes. A major issue with these datasets is that they contain images from the same scene in both the training and test sets, potentially affecting generalization.
Several datasets were subsequently introduced to address the limitations of early LLIE benchmarks, primarily in terms of scale and diversity. The DPED [11] dataset provided images captured across multiple smartphone cameras, enabling cross-device evaluation. Deep-UPE [30] emphasized extremely low-light scenarios with more challenging exposure conditions. The LSRW [9] dataset expanded camera diversity by including both DSLR and smartphone captures, recognizing the distinct image formation characteristics of different sensor types. More recently, LLIV-Phone [15] introduced temporal information by capturing video sequences under low-light conditions, allowing methods to exploit inter-frame correlations. This video-based approach was further explored in the DID [5] and SDSD [29] datasets, which provided paired low-light and normal-light video sequences for dynamic scene enhancement.
A significant methodological shift came with the SID [3] dataset, which captured image pairs in RAW format rather than processed RGB, enabling methods to leverage the complete sensor information before in-camera processing. To scale dataset creation, VE-LOL [17] adopted a synthetic approach, darkening well-exposed images and adding synthetic noise patterns to simulate sensor characteristics at high ISO settings. Recently, the BVI-LowLight dataset [19] was introduced, containing 40,000 images of objects captured at different ISO settings. Additionally, to obtain more training data, some LLIE methods use exposure-correction datasets such as PHOS [26] and SICE [1]. Despite these advances in scale, sensor diversity, temporal modeling, and data modality, existing datasets either capture each scene under a single low-light condition or rely on modifying the camera parameters.
@@ 2.2 Low-Light Image Enhancement Methods
Early LLIE methods built upon the seminal Retinex algorithm [13], which decomposed images into reflectance and illumination components, inspiring variants including Multi-Scale Retinex [22], SRIE [6], and Milano-Retinex [25]. Methods such as LIME [8] and NPE [31] demonstrated strong performance by leveraging natural image statistics without training data. However, these traditional approaches have been largely superseded by deep learning methods.
Early end-to-end deep learning methods include SID [3] for RAW images and RetinexNet [35] for RGB inputs. Subsequent approaches introduced various architectural innovations: GLADNet [34] combined global illumination and local detail modules; KinD [47] and KinD++ [46] adopted Retinex-inspired decomposition strategies; and Yang et al. [38] incorporated adversarial learning. Recent methods leverage transformer architectures (LLFormer [43, 32], Retinexformer [2]) and diffusion models (Diff-Retinex [40], PyDiff [48]). Alternative formulations include specialized color spaces [37] and pixel-wise mean estimation losses [16]. Some methods jointly address enhancement and degradation removal, such as DarkIR [4] and LEDNet [49]. To avoid paired training data requirements, unsupervised approaches have been proposed, including EnlightenGAN [12], Zero-DCE [7], SCI [18], and lightweight RUAS [23].
However, all existing methods have been evaluated exclusively on fixed single low-light inputs without considering behavior across varying illumination levels. We benchmark state-of-the-art LLIE methods across different brightness conditions and propose two simple modifications to Retinexformer [2] that leverage our multi-illumination dataset to improve robustness across intensity levels.
@@ 3 Multi-Illumination Low-Light (MILL) Dataset
Figure 3: Example scenes from our dataset at different levels for both the DSLR (first two rows) and the smartphone camera (last row).
Existing LLIE datasets present two critical limitations: (1) they either contain a single severely underexposed image per scene (e.g. [39, 35]) or (2) simulate brightness variations via camera parameter adjustments or post-processing (e.g. [19, 1, 26]). This constraint limits real-world applicability, where low-light conditions span a continuous range of intensities.
To quantify this limitation, we simulated varying brightness levels on the LoLv1 dataset [33] by blending input images with their ground truth counterparts at different ratios. This blending reduces degradation severity by simulating intermediate brightness levels, theoretically making enhancement easier. As shown in Table 1, both Retinexformer [2] and CIDNet [37] performed worse on the blended versions (with 0.2 and 0.5 ground truth mixing ratios) than on the original dataset, as measured by $\Delta E_{76}$ and PSNRL. This result occurs because models trained on fixed brightness levels fail to generalize across different intensities. Figure 2 illustrates this problem. Oversaturation in the output increases proportionally with input brightness, a clear evidence that intermediate brightness levels are absent from training data. This lack of intensity diversity severely limits the practical applicability of LLIE methods, as real-world deployment requires robustness across varying brightness conditions. To study this problem and address this limitation, we introduce a novel LLIE dataset featuring multiple brightness levels per scene with fixed camera parameters, enabling more robust training, benchmarking, and evaluation of LLIE methods across multiple intensity levels.
Our primary objective is to capture a high-quality, well-calibrated dataset for evaluating and training LLIE methods across different intensity levels. To ensure consistency and eliminate uncontrollable variables, we captured all images in a controlled indoor environment without windows or external light sources. We used a dedicated room with a platform for placing different floor backgrounds and objects, equipped with programmable lighting to precisely control brightness levels. Images were captured using two devices: a Nikon D5200 DSLR camera and a Samsung Galaxy S7 smartphone. The smartphone provides a contrasting capture profile compared to the DSLR, enabling evaluation across different sensor characteristics. We employed nine Philips Hue lights, whose emission spectrum covers the full visible range and provides sufficient spectral support for our experiments. Although we acknowledge that LED-based illuminants can exhibit spectral peaks and may not perfectly reproduce the continuous spectra of natural Planckian light sources, they are the preferred technology for fine-grained intensity control. Figure 1 provides a schematic overview.
@@
RUAS [23] 0.003 CVPR’21 25.46 16.63 45.33 9.86 57.47 6.93 62.99 5.86 67.14 5.18
LLFormer [32] 24.52 AAAI’23 16.37 20.88 13.73 21.79 13.34 22.06 13.17 22.25 12.90 22.55
KinD [47] 1.20 ACMM’19 23.88 16.71 17.62 21.79 14.87 21.34 14.49 21.63 15.00 21.36
FourLLIE [28] 0.12 ACMM’23 24.51 17.29 22.79 17.66 26.64 14.97 28.79 14.07 30.79 13.35
SCI [18] 0.0003 CVPR’22 24.05 16.02 17.99 21.42 25.89 15.69 31.66 13.24 38.38 11.23
@@ 4 Method using New Loss Terms
Most current LLIE architectures follow a UNet-like structure, comprising an encoder and a decoder. We aim to disentangle the latent features extracted by the architecture’s bottleneck. We adopt Retinexformer [2] as our baseline architecture due to its strong performance in our benchmark evaluation (see Section 5.1).
@@ 4.1 Intensity Prediction Loss
KinD [47] 20.409 18.157 0.779 0.219 16.782 1.871 4.383 15.948
FourLLIE [28] 15.486 13.534 0.687 0.241 26.759 3.514 5.398 19.201
SCI [18] 15.329 13.006 0.609 0.270 28.669 3.407 5.791 26.788
@@
Smartphone Unprocessed 19.155 17.573 0.511 0.215 17.832 2.718 5.088 20.895
RUAS [23] 6.920 5.082 0.267 0.722 62.707 6.246 10.556 71.760
@@
KinD [47] 19.462 17.452 0.536 0.259 17.252 1.980 3.546 22.950
FourLLIE [28] 21.377 18.510 0.539 0.236 16.976 3.279 4.831 22.159
SCI [18] 15.960 13.208 0.437 0.330 29.477 3.314 5.475 24.372
MirNet [41] 20.760 18.100 0.614 0.219 15.231 1.720 3.537 23.269
@@ 5.1 Benchmark on MILL
We benchmark mainstream LLIE methods by retraining them on our MILL-s dataset using their officially released code. Our evaluation includes unsupervised methods (RUAS [23], SCI [18]), a Retinex-based approach (KinD [47]), transformer-based methods built on Restormer [42] (LLFormer [43], Retinexformer [2], PromptNorm [24]), a frequency-domain method (FourLLIE [28]), image restoration approaches (MIRNet [41], DarkIR [4]), and specialized LLIE methods (HVI-CIDNet [37], GT-Mean [16]). We also evaluate our proposed modifications to Retinexformer.
Input SCI [18] GT-Mean [16] PromptNorm [24] Retinexformer [2] Ours Ground Truth
Figure 4: Visual comparison on MILL-s. From left to right: input, SCI [18], GT-Mean [16], PromptNorm [24], Retinexformer [2], Ours, and ground truth. We show three examples with zoomed-in regions below each. First two images: DSLR; last image: smartphone.
@@
Figure 6: Outdoor examples from the DICM [14] (first row) and SICE [1] (second row) of Retinexformer trained on LoLv1, our baseline with the two proposed additional loss terms independently, and our final approach.
@@
Table 3 reports performance averaged across all intensity levels for both the DSLR and smartphone splits of MILL-s, using three full-reference metrics: PSNRL, PSNRC (on RGB), and SSIM; one perceptual full-reference metric: LPIPS [44]; two color similarity metrics: $\Delta E_{76}$ and MS-SWD [10]; and two non-reference metrics: NIQE [21], and Brisque [20]. Surprisingly, several methods fail to improve upon the input images on average. As demonstrated in Table 5, these methods enhance extremely low-light images but degrade moderately underexposed images, resulting in net negative impact. This observation highlights the difficulty of achieving robust LLIE across variable intensity levels. Retinexformer achieves the best performance, followed closely by PromptNorm and LLFormer. This indicates that Restormer-based architectures perform well for this task. We therefore select Retinexformer as our baseline due to its superior performance and parameter efficiency. Our proposed method further improves upon Retinexformer. The smartphone split proves more challenging across all methods due to inferior sensor quality.
@@ 5.3 FullHD Experiments and Ablation
Our proposed modifications outperform the baseline model across all metrics. On the DSLR split, we observe improvements of approximately 10 dB in PSNRL and PSNRC, 0.03 in SSIM, and 5 in $\Delta E{76}$. The smartphone split exhibits smaller but consistent gains due to sensor limitations; nevertheless, our method maintains a clear performance advantage over the baseline. Notably, the intensity prediction loss yields larger improvements than the scene content loss when applied independently. However, combining both losses delivers the strongest performance, as effective feature disentanglement requires their joint optimization. Given the performance gap with respect to Retinexformer, we applied a global intensity adjustment to the baseline model; however, the results changed minimally, confirming that the improvements stem from feature disentanglement rather than brightness correction.